# Training objective drives the consistency of representational similarity across datasets

### v1

@@ 4 Experiments
Our experiments investigate the consistency of representational similarities between vision models across various datasets using our proposed framework. Following the experimental setup (Sec. 4.1), we first analyze how model similarities transfer across datasets (Sec. 4.2 and 4.4) and find consistent model groups through similarity clustering (Sec. 4.3). We then investigate how model and dataset characteristics influence similarity consistency (Sec. 4.5 and 4.6), and conclude by examining the relationship between representational similarities and models’ performance differences on downstream tasks (Sec. 4.7). All code to run our analyses and reproduce the results presented in this section is publicly available on GitHub: https://github.com/lciernik/similarity_consistency.
@@ 4.1 Setup
Models. We evaluated 64 vision models, representing a diverse range of architectures, training paradigms, and parameter scales. A complete list of models and their characteristics can be found in Appx. A. We used the Python package thingsvision [62] to extract the model representations.

### v2

@@ 4 Experiments
Our experiments investigate the consistency of representational similarities between vision models across various datasets using our framework. Following the experimental setup (§4.1), we first analyze how model similarities transfer across datasets (§4.24.4) and find consistent model groups through similarity clustering (§4.3). We then investigate how model and dataset characteristics influence similarity consistency (§4.5 and 4.6), and conclude by examining the relationship between representational similarities and models’ performance differences on downstream tasks (§4.7). The code and the data to run our analyses and reproduce the experimental results are publicly available at https://github.com/lciernik/similarity_consistency.
@@ 4.1 Setup
Models. We evaluated 64 vision models, representing a diverse range of architectures, training paradigms, and parameter scales. We focused on general-purpose vision models trained on large-scale datasets (minimum ImageNet-1k size) with broad semantic diversity, excluding models trained on specialized or synthetic datasets. A complete list of models and their characteristics can be found in Appx. A.